# Concurrent-Learning Based Relative Localization in Shape Formation of Robot Swarms

### v1

@@ V-B2 Finite time consensus based estimator design
$+\mu_{i}(\hat{q}_{i,0}(t)-\hat{p}_{i0,0}(t))\Big{)}^{\alpha},$ (17)
where $c_{1}>0$ and $10<\alpha<$ 1 are two positive constants. Parameter $a_{ij}=1$ means that robot $j$ is a neighbor of robot $i$ at initial time $t_{0}$, otherwise $a_{ij}=0$. Parameter $\mu_{i}=1$ means that robot $i$ is a neighbor of the seed robot, otherwise $\mu_{i}=0$. For vector $x=[x_{1},...,x_{n}]^{T}$, function $\mathrm{sig}(x)^{\alpha}=\sum_{i=1}^{n}\mathrm{sign}(x_{i})\lvert x_{i}\rvert^{\alpha}$, where $\mathrm{sign}(\cdot)$ is the sign function. Recall that the relative position estimator $\hat{p}_{ij,0}$ defined in (7) is also adopted in (V-B2). Define the estimated error as $\tilde{q}_{i,0}=\hat{q}_{i,0}-p_{i0}(t_{0})$, take $\tilde{q}=[\tilde{q}_{1,0},...,\tilde{q}_{n,0}]^{T}$.
@@
Note that the connected constrain of $\mathcal{G}(t_{0})$ in Assumption 2 is common and can be found in several consensus results including [43, 44, 45]. The constrain of $r_{\mathrm{neigh}}$ is to guarantee that the initial neighbor robots in set $\mathcal{N}_{i}(t_{0})$ are still the neighbors of robot $i$ during the shape agreement process. Next, we will show the convergence of the agreement protocol (V-B2) in the following theorem. To achieve this, we choose a Lyapunov function candidate as,
$V_{l}=\frac{1}{2}\tilde{q}^{T}(\mathcal{L}+\mathcal{B})\tilde{q},$ (19)
where $\mathcal{B}=\mathrm{diag}(\mu_{1},...,\mu_{n})$. Let $\mathcal{L}$ be the Laplace matrix of the initial measurement topology $\mathcal{G}(t_{0})$.
@@ 27            $\hat{q}_{i,0}$leftarrow update shape position estimator according to (V-B2);
$b=\sqrt{\frac{\frac{\alpha}{1+\alpha}2^{\frac{1+3\alpha}{2\alpha(\alpha+1)}}\left(\sum_{i=1}^{n}(|\mathcal{N}_{i}(t_{0})|\varepsilon)^{1+\alpha}\right)}{(1-\gamma)\frac{\alpha}{1+\alpha}\lambda_{\min}(\mathcal{L}+\mathcal{B})^{2}}},$ (20)
$t_{l}=\frac{2(1+\alpha)V_{l}(t_{a})^{\frac{1-\alpha}{2}}}{\lambda_{\min}(\mathcal{L}+\mathcal{B})\gamma(1-\alpha)},$ (21)
$t_{a}=\mathrm{\max}\{\frac{\mathrm{ln}(\varepsilon\lVert\tilde{p}_{ij,0}(t_{0})\rVert^{-1})}{\mathrm{ln}(\lambda_{ij})}\Delta t,(i,j)\in\mathcal{G}(t_{0})\},$ (22)
@@
However, calculating convergence time requires the initial measurement topology of the swarm, which is unavailable in real applications. To overcome this issue, a gradient threshold based decision method and the hop-count information transmission algorithm [11] are adopted in this paper. Function $\mathrm{CalUpdRate(\cdot)}$ calculates the variation rate of $\hat{q}_{i,0}$ over a period of time $\delta_{t}$ which can be defined as,
$\Delta\hat{q}_{i,0}=\frac{\hat{q}_{i,0}(t)-\hat{q}_{i,0}(t-\delta_{t})}{\delta_{t}}.$ (23)
If $\Delta\hat{q}_{i,0}$ defined as (23) remains above a user-defined threshold convergence rate $\delta_{0}$, the hop-count $hop_{i}$ is set as 0. Otherwise, it means that the robots have reached an approxiate agreement, the hop-count $hop_{i}$ is set as $\mathrm{\min}(hop_{j})+1$, where $j\in\mathcal{N}_{i}(t_{0})$ (Algorithm 1, Lines 18-22). The value of hop-count increases rapidly when all the robots reach an agreement.
@@ V-C Distributed shape formation
The proposed behavior based control strategy consists of two stages, i.e., the shape localization agreement (Algorithm 1, line 9-17) and shape formation (Algorithm 2). During the shape localization agreement stage, if robot $i$ does not have unlocalized neighboring robots, the behavior based control law is designed as
@@
When all the robots reach an consensus on the shape localization with the agreement protocol (V-B2). The shape formation stage begins. During this stage, if robot $i$ does not have unlocalized neighboring robots, the control law is designed as
@@
Here, $v_{i}^{\mathrm{ent}}$, $v_{i}^{\mathrm{exp}}$ and $v_{i}^{\mathrm{int}}$ stand for the shape-entering command, shape-exploring command and neighbor-interacting command, respectively.
@@ 8             $\mathcal{RL}_{i}$leftarrow update the localized robot set;
10             $v_{i}^{\mathrm{ent}}$leftarrow calcaulate shape forming command (28);
11             $v_{i}^{\mathrm{exp}}$leftarrow calcaulate shape exploration command (29);
12             $v_{i}^{\mathrm{int}}$leftarrow calcaulate interation command (30);
@@ 18             $\mathcal{RD}_{i}$leftarrow record current measurement $\mathcal{CD}_{i}$;
19             $v_{i}^{\mathrm{enh}}$leftarrow calcaulate localization enhance command (31);
20             $v_{i}^{\mathrm{int}}$leftarrow calcaulate interation command (30);
@@ 26             $\hat{p}_{ij,0}$leftarrow update relative position esetimator according to (7)
According to the control law designed as (24) and (25), one can see that each robot either makes circular localization-enhancing motion or stops during the shape localization agreement stage. As a result, the collision avoidance between robots can be easily guaranteed by adjusting the initial distance between the two neighboring robots. During the shape formation stage, the collision avoidance between robots can be achieved by the neighbor-interacting command $v_{i}^{\mathrm{int}}$ in (26) and (27). The design of the four control commands are as follows.
@@ V-C1 Shape-entering command
When all the neighboring robots of robot $i$ are localized, robot $i$ turns to form the desired shape (Algorithm 2, lines 4-11). The shape-entering command $v_{i}^{\mathrm{ent}}$ is designed to derive robot $i$ to its interpretation of the desired shape, as depicted in Fig. 2C. In particular, it is designed as
$v_{i}^{\mathrm{ent}}=\kappa_{1}\xi_{\rho_{T,i}}\frac{\hat{q}_{i}+p_{\rho_{T,i}0}}{\hat{q}_{i}+p_{\rho_{T,i}0}},$ (28)
where $\kappa_{1}>0$ is a constant. Recall that $\hat{q}_{i}$ is the real-time shape localization estimate (18). The relative position between the nearest gray cell to robot $i$ is denoted as $p_{\rho_{T,i}0}$. Note that $p_{\rho_{T,i}0}$ can be determined by $\hat{q}_{i}$ and the gray formation shape $\mathcal{F}_{\mathrm{gray}}$. The gray level of the cell is $\xi_{\rho_{T,i}}$.
@@ V-C2 Shape-exploring command
$v_{i}^{\mathrm{exp}}=\frac{\sum_{\rho\in\mathcal{M}_{i}^{\mathrm{sense}}}\kappa_{2}\phi(\lVert\hat{q}_{i}+p_{\rho 0}\rVert)\left(\hat{q}_{i}+p_{\rho 0}\right)}{\sum_{\rho\in\mathcal{M}_{i}^{\mathrm{sense}}}\phi(\lVert\hat{q}_{i}+p_{\rho 0}\rVert)},$ (29)
where $\kappa_{2}>0$ is a constant. Function $\phi(z)$ is defined as,
$\phi(z)=\left\{\begin{aligned} &1&z\leq 0\\ &\frac{1}{2}(1+\mbox{cos}\pi z)\quad&0<z<1\\ &0&z\geq 1\end{aligned},\right.\end{aligned}$
@@ V-C3 Neighbor-interacting command
As shown in Fig. 2C, the interaction command $v_{i}^{\mathrm{int}}$ is to achieve collision avoidance and velocity alignment between robots. To this end, it is designed as
$v_{i}^{\mathrm{int}}=\kappa_{3}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{RL}_{i}}\mu(d_{ij})\hat{p}_{ij}+\kappa_{4}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{RL}_{i}}(v_{i}-v_{j}),$ (30)
where $\kappa_{3}>0$ and $\kappa_{4}>0$ are two constants. Here, $\mu(z)$ is defined as
@@ V-C4 Localization-enhancing command
During the process of shape formation control, when robot $i$ encounters a neighboring robot $j$ which is not localized, i.e., $j\in\mathcal{N}_{i}$ and $j\notin\mathcal{RL}_{i}$. It makes localization-enhancing motion and collects measurement data to achieve localization(Algorithm 2, Lines 12-18).
According to the theoretical analysis in Section IV-D, the localization-enhancing command $v_{i}^{\mathrm{enh}}$ is designed as
$v_{i}^{\mathrm{enh}}=[r^{{}^{\prime}}_{i}w_{i}\mathrm{cos}(\varphi),r^{{}^{\prime}}_{i}w_{i}\mathrm{sin}(\varphi)]^{T},$ (31)
@@
Compared with our previous work [8], we focus on the shape formation control with only local measurements in this paper. To achieve this goal, a localization-enhancing distributed control strategy is proposed. All the control commands are based on the estimated relative position. The relative position estimator (7) with exponential convergence and the center negotiation estimator (V-B2) with finite time convergence play an important role in the close-loop control strategy.
@@ VI Simulation Results and Performance Analysis
In this section, simulation results are provided to demonstrate the robustness, effectiveness and adaptability of the proposed strategy.
Figure 4: Comparison between the proposed relative localization method and the PE based one in [18]. A. Without measurement noise, numerical simulation results with the proposed method. From left to right are the trajectory and tracking error of each robot. Here, the tracking error is defined as $e_{i}^{\mathrm{c}}(t)=\lVert p_{i}(t)-p_{0}-p_{i}^{*}\rVert$. B. Without measurement noise, numerical simulation results with the method in [18]. From left to right are the trajectory and tracking error of each robot. C. With measurement noise, numerical simulation results with the proposed method. From left to right are the trajectory, tracking error, estimation error and velocity norm of each robot. Here, the estimation error is defined as $e_{i}^{\mathrm{e}}(t)=\lVert\hat{p}_{i,0}(t)-p_{0}-p_{i}\rVert$. D. With measurement noise, numerical simulation results with the method in [18]. From left to right are the trajectory, tracking error, estimation error and velocity norm of each robot.
@@ VI-A Performance of Relative Localization Method
First, we evaluate the performance of the proposed relative localization method by a distributed formation scenario.
In the case without measurement noise, five robots are assigned to achieve formation with the proposed concurrent-learning based relative localization method and the localization method proposed in [18]. The position of a seed robot is chosen as $p_{0}=[0,7]^{T}$ and the reference formation offsets for another four robots are chosen as $p*=[p_{1}^{*T},p_{2}^{*T},p_{3}^{*T},p_{4}^{*T}]^{T}=[-7,-7,0,-7,0,-14,7,-7]^{T}$. The measurement topology is fixed, the connectivity matrix $\mathcal{A}$ and informed matrix $\mathcal{B}$ are $\mathcal{A}=[0,1,0,0;1,0,1,0;0,1,0,1;0,0,1,0]$ and $\mathcal{B}=\mathrm{diag}(1,1,0,0)$. The control input of each robot is simply chosen as the following P-controller,
$v_{i}(t)=-\kappa(\hat{q}_{i}(t)-p_{i}^{*}),$
where $\kappa=0.02$. Parameters adopted in the relative position estimator are $\Delta t=0.01$ $\lambda_{0}=0.1$, $c_{1}=0.005$, $\alpha=0.1$, and the data batch size is chosen as $h=60$. Due to the fixed measurement topology, the robots can move and estimate the relative position of the seed robot at the same time. As a result, there is no need for parameter $\delta_{t}$ and $\delta_{0}$ in this scenario. Parameters adopted in the contrast simulation can be found in [18]. The trajectories of the robots and the tracking errors are shown in Fig. 4A and 4B. It can be seen that without the persistent noise, the trajectories of the robots become smoother in Fig. 4A.
Furthermore, to verify the noise robustness of the relative localization algorithm, white Gaussian noise with zero mean and variance of $0.02$m are added to distance measurement. Besides, the zero mean Gaussian noise with a variance of $0.002$m/s is also added to the odometry measurement. Parameters adopted in this case are the same with those in the noise-free case. The trajectories of the robots, the norm of tracking errors, the norm of estimation errors and the norm of robot velocities are shown in Fig. 4C and 4D.
From the simulation results, it can be seen that with the measurement noise, the performance of both the two localization methods deteriorate inevitably. However, compared with method proposed in [18], the concurrent-learning based estimation algorithm has better robustness against noisy measurements. This is due to the fact that our method simultaneously utilizes historical data for information enhancement when estimating the relative position.

### v4

@@ V-B2 Finite time consensus based estimator design
$+\mu_{i}(\hat{q}_{i,0}(t)-\hat{p}_{i0,0}(t))\Big{)}^{\alpha}$ (14)
where $c_{1}>0$ and $10<\alpha<$ 1 are the gain and exponent of the estimator (V-B2). Parameter $a_{ij}=1$ means that robot $j$ is a neighbor of robot $i$ at initial time $t_{0}$, otherwise $a_{ij}=0$. Parameter $\mu_{i}=1$ means that robot $i$ is a neighbor of the seed robot, otherwise $\mu_{i}=0$. Define $\mathcal{B}=\mathrm{diag}(\mu_{1},...,\mu_{n})$. For vector $x=[x_{1},...,x_{n}]^{T}$, function $\mathrm{sig}(x)^{\alpha}=\sum_{i=1}^{n}\mathrm{sign}(x_{i})\lvert x_{i}\rvert^{\alpha}$, where $\mathrm{sign}(\cdot)$ is the sign function. Recall that the relative position estimator $\hat{p}_{ij,0}$ defined in (6) is also adopted in (V-B2). Define the estimated error as $\tilde{q}_{i,0}=\hat{q}_{i,0}-p_{i0}(t_{0})$, take $\tilde{q}=[\tilde{q}_{1,0},...,\tilde{q}_{n,0}]^{T}$.
@@
Note that the connected constrain of $\mathcal{G}(t_{0})$ in Assumption 2 is common and can be found in several consensus results including [13, 8]. The constrain of $r_{\mathrm{neigh}}$ is to guarantee that the initial neighbor robots in set $\mathcal{N}_{i}(t_{0})$ are still the neighbors of robot $i$ during the shape agreement process. Let $\mathcal{L}$ be the Laplace matrix of the initial measurement topology $\mathcal{G}(t_{0})$. The convergence of the agreement protocol (V-B2) can be given in the following theorem. Choose a Lyapunov function candidate,
$V_{l}=\frac{1}{2}\tilde{q}^{T}(\mathcal{L}+\mathcal{B})\tilde{q}$ (16)
@@ 27            $\hat{q}_{i,0}$leftarrow update shape position estimator according to (V-B2);
$b=\sqrt{\frac{2^{\frac{1+3\alpha}{2\alpha(\alpha+1)}}\left(\sum_{i=1}^{n}(|\mathcal{N}_{i}(t_{0})|\varepsilon)^{1+\alpha}\right)}{(1-\gamma)\lambda_{\min}(\mathcal{L}+\mathcal{B})^{2}}}$ (17)
$t_{l}=\frac{2(1+\alpha)V_{l}(t_{a})^{\frac{1-\alpha}{2}}}{\lambda_{\min}(\mathcal{L}+\mathcal{B})\gamma(1-\alpha)}$ (18)
$t_{a}=\mathrm{\max}\{\frac{\mathrm{ln}(\varepsilon\lVert\tilde{p}_{ij,0}(t_{0})\rVert^{-1})}{\mathrm{ln}(\lambda_{ij})}\Delta t,(i,j)\in\mathcal{G}(t_{0})\}$ (19)
@@
However, calculating convergence time requires the initial measurement topology of the swarm, which is unavailable in real applications. To overcome this issue, a gradient threshold based decision method and the hop-count information transmission algorithm [7] are adopted in this article. Function $\mathrm{CalUpdRate(\cdot)}$ calculates the variation rate of $\hat{q}_{i,0}$ over a period of time $\delta_{t}$ which can be defined as,
$\Delta\hat{q}_{i,0}=\frac{\lVert\hat{q}_{i,0}(t)-\hat{q}_{i,0}(t-\delta_{t})\rVert}{\delta_{t}}.$ (20)
If $\Delta\hat{q}_{i,0}$ defined as (20) remains above a user-defined threshold convergence rate $\delta_{0}$, the hop-count $hop_{i}$ is set as 0. Otherwise, it means that the robots have reached an approximate agreement, the hop-count $hop_{i}$ is set as $\mathrm{\min}(hop_{j})+1$, where $j\in\mathcal{N}_{i}(t_{0})$ (Algorithm 1, Lines 18-22). The value of hop-count increases rapidly when all robots reach an agreement.
@@ V-C Distributed Shape Formation
The proposed behavior based control strategy consists of two stages, i.e., the shape localization agreement (Algorithm 1, lines 9-17) and shape formation (Algorithm 2). During the shape localization agreement stage, if robot $i$ does not have unlocalized neighbors, the control law $v_{i}$ is designed as
@@
When all the robots reach a consensus on the shape localization with the agreement protocol (V-B2). The shape formation stage begins. During this stage, if robot $i$ does not have unlocalized neighboring robots, $v_{i}$ is designed as
@@
Here, $v_{i}^{\mathrm{ent}}$, $v_{i}^{\mathrm{exp}}$ and $v_{i}^{\mathrm{int}}$ denote the shape-entering command, shape-exploring command and neighbor-interacting command.
@@ 8             $\mathcal{RL}_{i}$leftarrow update the localized robot set;
10             $v_{i}^{\mathrm{ent}}$leftarrow calculate shape forming command (25);
11             $v_{i}^{\mathrm{exp}}$leftarrow calculate shape exploration command (26);
12             $v_{i}^{\mathrm{int}}$leftarrow calculate interaction command (27);
@@ 18             $\mathcal{RD}_{i}$leftarrow record current measurement $\mathcal{CD}_{i}$;
19             $v_{i}^{\mathrm{enh}}$leftarrow calculate localization enhance command (28);
20             $v_{i}^{\mathrm{int}}$leftarrow calculate interaction command (27);
@@ 26             $\hat{p}_{ij,0}$leftarrow update relative position esetimator according to (6)
According to the control law designed as (21) and (22), one can see that each robot either performs circular localization-enhancing motion or stops during the shape localization agreement stage. As a result, the collision avoidance between robots can be achieved if the initial distance $d_{ij}(t_{0})$ between them satisfies $d_{ij}(t_{0})>2(r_{i}+r_{j})$. During the shape formation stage, the collision avoidance between robots can be achieved by the neighbor-interacting command $v_{i}^{\mathrm{int}}$ in (23) and (24). The four control commands are designed as follows.
@@ V-C1 Shape-entering command
When all the neighboring robots of robot $i$ are localized, robot $i$ turns to form the desired shape (Algorithm 2, lines 4-11). The shape-entering command $v_{i}^{\mathrm{ent}}$ is designed to derive robot $i$ to its interpretation of the desired shape, as depicted in Fig. 2C. The shape-entering command can also be understood as a proportional controller, driving the robots into the shape.
It is designed as
$v_{i}^{\mathrm{ent}}=\kappa_{1}\xi_{\rho_{T,i}}\frac{\hat{q}_{i}+p_{\rho_{T,i}0}}{\hat{q}_{i}+p_{\rho_{T,i}0}}$ (25)
where $\kappa_{1}>0$ is the gain of the shape-entering command.The larger $\kappa_{1}$, the greater the force that pushes the robot into shape. Recall that $\hat{q}_{i}$ is the real-time shape localization estimate (15). The relative position between the nearest gray cell to robot $i$ is denoted as $p_{\rho_{T,i}0}$. Note that $p_{\rho_{T,i}0}$ can be determined by $\hat{q}_{i}$ and the gray formation shape $\mathcal{F}_{\mathrm{gray}}$. The gray level of the cell is $\xi_{\rho_{T,i}}$.
@@ V-C2 Shape-exploring command
$v_{i}^{\mathrm{exp}}=\frac{\sum_{\rho\in\mathcal{M}_{i}^{\mathrm{sense}}}\kappa_{2}\phi(\lVert\hat{q}_{i}+p_{\rho 0}\rVert)\left(\hat{q}_{i}+p_{\rho 0}\right)}{\sum_{\rho\in\mathcal{M}_{i}^{\mathrm{sense}}}\phi(\lVert\hat{q}_{i}+p_{\rho 0}\rVert)}$ (26)
where $\kappa_{1}>0$ is the gain of the shape-exploring command. The larger $\kappa_{2}$, the greater the force exerted by the robot in shape exploration. Function $\phi(z)$ is defined as,
$\phi(z)=\left\{\begin{aligned} &1&z\leq 0\\ &\frac{1}{2}(1+\mbox{cos}\pi z)\quad&0<z<1\\ &0&z\geq 1\end{aligned}\right.\end{aligned}$
@@ V-C3 Neighbor-interacting command
As shown in Fig. 2C, the interaction command $v_{i}^{\mathrm{int}}$ is to achieve collision avoidance and velocity alignment between robots, it is designed as
$v_{i}^{\mathrm{int}}=\kappa_{3}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{RL}_{i}}\mu(d_{ij})\hat{p}_{ij}+\kappa_{4}\sum_{j\in\mathcal{N}_{i}\cap\mathcal{RL}_{i}}(v_{i}-v_{j})$ (27)
where $\kappa_{3}>0$ and $\kappa_{4}>0$ are the gains of collision-avoidance and velocity-alignment, respectively. The larger $\kappa_{3}$, the greater the collision avoidance force between robots, similar to $\kappa_{4}$. $\mu(z)$ is defined as
@@ V-C4 Localization-enhancing command
During the process of shape formation control, when robot $i$ encounters a neighboring robot $j$ which is not localized, i.e., $j\in\mathcal{N}_{i}$ and $j\notin\mathcal{RL}_{i}$. $v_{i}^{\mathrm{enh}}$ drive it to make localization-enhancing motion and collect measurement data to achieve localization(Algorithm 2, Lines 12-18). According to the theoretical analysis in Section IV-D, localization-enhancing command $v_{i}^{\mathrm{enh}}$ is designed as
$v_{i}^{\mathrm{enh}}$ $=[r^{{}^{\prime}}_{i}w_{i}\mathrm{cos}(\varphi),r^{{}^{\prime}}_{i}w_{i}\mathrm{sin}(\varphi)]^{T}$ (28)
@@
Figure 4: Comparision between the proposed relative localization method and the PE based one in [14]. A. Simulation results with the proposed method, from top to bottom are the trajectory, estimate error, tracking error and velocity norm of each robot. Here, tracking error and estimation error are defined as $e_{i}^{\mathrm{c}}(t)=\lVert{p}_{i}(t)-p_{0}-p_{i}\rVert$ and $e_{i}^{\mathrm{e}}(t)=\lVert\hat{p}_{i,0}(t)-p_{0}-p_{i}\rVert$, respectively. B. Simulation results with method in [14].
@@
Compared with our previous work [5], we focus on the shape formation control with only local measurements in this article. To achieve this goal, a localization-enhancing distributed control strategy is proposed. All the control commands are based on the estimated relative position. The relative position estimator (6) with exponential convergence and the center negotiation estimator (V-B2) with finite time convergence play an important role in the close-loop control strategy.
@@ VI Simulation Results and Performance Analysis
Simulation results are given in this section to show the robustness, effectiveness and adaptability of our methods.
Figure 5: Comparision between the proposed relative localization method and the PE based one in [14] with the sensor measurement noise. A. Simulation results with the proposed method, from top to bottom are the trajectory, estimate error, tracking error and velocity norm of each robot. B. Simulation results with method in [14].
@@ VI-A Performance of Relative Localization Method
First, we evaluate the performance of the proposed relative localization method by a distributed formation scenario. In the case without measurement noise, five robots are assigned to achieve formation with the proposed concurrent-learning based relative localization method and the localization method proposed in [14]. The position of a seed robot is chosen as $p_{0}=[0,7]^{T}$ and the reference formation offsets for another four robots are chosen as $p*=[p_{1}^{*T},p_{2}^{*T},p_{3}^{*T},p_{4}^{*T}]^{T}=[-7,-7,0,-7,0,-14,7,-7]^{T}$. The measurement topology is fixed, the connectivity matrix $\mathcal{A}$ and informed matrix $\mathcal{B}$ are $\mathcal{A}=[0,1,0,0;1,0,1,0;0,1,0,1;0,0,1,0]$ and $\mathcal{B}=\mathrm{diag}(1,1,0,0)$. The control input of each robot is $v_{i}(t)=-\kappa(\hat{q}_{i}(t)-p_{i}^{*})$, where $\kappa=0.2$. The localization enhancement motion radius and angle velocity are $r_{i}=0.6$m, $v_{\mathrm{max}}=0.75$m/s, $w_{0}=1$rad/s, $w_{r}=1$rad/s. The parameters adopted in the relative position estimator are $\Delta t=0.01$ $\lambda_{0}=0.1$, $c_{1}=0.005$, $\alpha=0.1$, and the data batch size is chosen as $h=60$. Note that the controller here is not the focus of our simulation. We have chosen the simplest P-controller, and other controllers such as PID or MPC can also be used in practical use to improve tracking control performance. Under the fixed communication and measurement topology, there is no need for parameters $\delta_{t}$ and $\delta_{0}$ in this scenario. The parameters adopted in the comparative simulation can be found in [14]. The trajectories of the robots and the tracking errors are shown in Fig. 4A and 4B. It can be seen that without the persistent noise, the trajectories of the robots become smoother and velocity command converges to zero quickly in Fig. 4A.
Furthermore, to verify the noise robustness of the relative localization algorithm, white Gaussian noises with zero mean and variance of $0.02$m and $0.002$m/s are added to distance and odometry measurement, respectively. The parameters adopted in this case are the same as those in the noise-free case. The trajectories of the robots, the norm of tracking errors, the norm of estimation errors and the norm of robot velocities are shown in Fig. 5C and 5D. From the simulation results, it can be seen that compared with the method proposed in [14], the concurrent-learning based estimation algorithm has better robustness against noisy measurements. This is due to the fact that our method simultaneously utilizes historical data for information enhancement.